# Calibrating Long-form Generations from Large Language Models

### v1

@@ Abstract
To enhance Large Language Models’ (LLMs) reliability, calibration is essential—the model’s assessed confidence scores should align with the actual likelihood of its responses being correct. However, current confidence elicitation methods and calibration metrics typically rely on a binary true/false assessment of response correctness. This approach does not apply to long-form generation, where an answer can be partially correct. Addressing this gap, we introduce a unified calibration framework, in which both the correctness of the LLMs’ responses and their associated confidence levels are treated as distributions across a range of scores. Within this framework, we develop three metrics to precisely evaluate LLM calibration and further propose two confidence elicitation methods based on self-consistency and self-evaluation. Our experiments, which include long-form QA and summarization tasks, demonstrate that larger models don’t necessarily guarantee better calibration, that calibration performance is found to be metric-dependent, and that self-consistency methods excel in factoid datasets. We also find that calibration can be enhanced through techniques such as fine-tuning, integrating relevant source documents, scaling the temperature, and combining self-consistency with self-evaluation. Lastly, we showcase a practical application of our system: selecting and cascading open-source models and ChatGPT to optimize correctness given a limited API budget. This research not only challenges existing notions of LLM calibration but also offers practical methodologies for improving trustworthiness in long-form generation.
@@
Figure 2: Overview of our calibration framework. First, we prompt an LLM to produce an answer to a specific question. Next, we assess the answer’s correctness distribution using an evaluator (GPT-4 or a task-specific metric). Subsequently, we determine the model’s confidence distribution through self-evaluation or self-consistency approaches. Finally, we calculate the calibration score by comparing the correctness and confidence distributions against our predefined metrics.
@@ 1 Introduction
Large language models (LLMs) are transforming the landscape of AI, yet their tendency to hallucinate—producing factually incoherent or unrealistic responses—poses a significant hurdle for real-world applications (Ji et al., 2022, Zhang et al., 2023b, Maynez et al., 2020, Guo et al., 2023, Sun et al., 2023, Min et al., 2023). One approach to harnessing their potential lies in calibration: augmenting predictions with confidence scores which are aligned with the likelihood of the predictions being correct. Correctness, however, is often subjective and granular, especially for long-form generation tasks like open-ended QA and summarization (Figure 1). For example, in an ambiguous question, an answer that only considers one of the possible interpretations is neither wrong nor totally correct (Stelmakh et al., 2022).
Current confidence estimation methods, on the other hand, predominantly focus on a binary notion of answer correctness and provide confidence scores for the whole sequence which reflect the statement “I am $x$confident that this answer is completely correct”. Such statements don’t necessarily apply to long-form generation, where correctness varies along a continuum. Furthermore, different annotators can assign varied correctness scores to the same answer Bakker et al. (2022), reflecting the ambiguity and subjectivity inherent in these tasks and their criteria. Given this, it becomes clear that representing an answer’s correctness as a single, definitive score is also limiting.
Addressing these challenges, we introduce a general calibration framework that conceptualizes the correctness of model responses as a distribution across scores between $[0,1]$ (Figure 1). We argue that most long-form generations lie somewhere in between, neither completely incorrect (00) nor perfect ($1$), and that representing the correctness as a distribution better allows capturing the subjectivity of human annotations or the variance in automated procedures for estimating correctness. Consequently, we view the confidence scores associated with the various correctness levels as another distribution over the same range $[0,1]$, corresponding to statements of the form “I am $x$confident that this answer is $y$correct”. Calibration is then defined as the degree of alignment between the predicted confidence distribution and the ground-truth correctness distribution obtained by evaluating the answers.
Methodologically, our framework encompasses three core components (Figure 2): estimating the ground-truth correctness distributions, eliciting confidence distributions from LLMs, and measuring calibration between these distributions. We introduce novel methods for each of these components: estimating correctness via prompting LLMs, eliciting confidence of long-form answers by adapting self-evaluation (Chiang and yi Lee, 2023) and self-consistency (Kadavath et al., 2022), and calibration metrics which measure both the divergence between distributions and their efficacy in selecting high-quality answers.
@@
We leverage our framework to measure calibration for several LLMs on multiple datasets across three long-form QA–including ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022)–and one summarization task, CNNDM (Nallapati et al., 2016). Our results show that stronger LLMs like GPT-3.5 don’t necessarily guarantee better calibration, calibration is metric-dependent, and LLMs exhibit better calibration with self-consistency method on factoid datasets than more open-ended datasets. Furthermore, our analysis reveals effective calibration enhancement strategies for long-form generation tasks. Fine-tuning and temperature scaling boost calibration across all metrics, while integrating relevant source documents and combining self-consistency with self-evaluation yield improvements only in part of metrics. Finally, we demonstrate a practical application of our framework: a cost-effective cascading strategy for using LLMs (Chen et al., 2023a). In this approach, open-source models initially handle queries and, based on their confidence levels—assessed using our system—more advanced API models are engaged as needed. This method ensures cost efficiency while maintaining high performance levels.
In essence, our contributions set a new benchmark for evaluating and improving LLM reliability across text generation tasks, offering:
  1. 1.
@@
    A comprehensive calibration framework applicable to all text generation tasks, enhancing LLM evaluation for critical applications.
  2. 2.
@@
  3. 3.
    Evidence that calibration can be augmented through model fine-tuning, document incorporation, temperature scaling, and mixed elicitation methods.
  4. 4.
    A cost-effective model usage strategy, illustrating the practicality of our calibration system in optimizing LLM deployment.
@@ 2 Related Work
Measuring Calibration Calibration (Guo et al., 2017, Minderer et al., 2021), the alignment of a model’s confidence with the actual accuracy of its predictions, has been widely studied in language models, whose probabilities derived from logits are generally found to not be calibrated (Jiang et al., 2020, Kadavath et al., 2022, Chen et al., 2023d). Standard metrics to measure the calibration include Expected Calibration Error (ECE) for confidence-accuracy disparity (Naeini et al., 2015), Brier Score for mean squared prediction-outcome differences, and AUROC for assessing confidence-based correct answer identification (Boyd et al., 2013, Kuhn et al., 2023). Selective Accuracy@Coverage measures accuracy within the model’s most confident predictions (Liang et al., 2023, Cole et al., 2023). However, these metrics are predominantly based on a binary conception of correctness, rendering them less effective for long-form text generation tasks where correctness is better represented as a distribution over a range of scores.
Improving Calibration Several studies develop improved methods for estimating confidence levels in model responses to enhance the calibration of LMs, such as training external calibrators (Mielke et al., 2022), fine-tuning models to express confidence (Lin et al., 2022), and using sensitivity to in-context learning as a proxy for confidence (Chen et al., 2023c). Given the rise of proprietary LLMs with limited access, confidence elicitation methods that don’t require access to logits (Hu et al., 2023) or embeddings have been extensively studied. These methods include verbalization-based, which directly asks the model to output its confidence (Lin et al., 2022), consistency-based, which samples multiple responses and uses consistency as a surrogate for confidence (Kadavath et al., 2022, Kuhn et al., 2023, Cole et al., 2023, Tian et al., 2023a), and the hybrid of both (Xiong et al., 2023, Tian et al., 2023b, Chen and Mueller, 2023). However, most existing methods still operate under the binary assumption of answer correctness. Consequently, confidence is a score that represents the model’s belief that the answer is all correct, inadequate for the complexity of long-form generation tasks.
3 Calibration Formulation
In the following sections, we first formalize the long-form generation calibration problem (§ 3.1), and then we introduce three core components in our calibration framework (Figure 2): the correctness distributions of the answers (§ 3.2), the confidence distributions of the LLM on its answers (§ 4), and the calibration metric to measure the difference between correctness distributions and confidence distributions (§ 5).
@@ 3.1 Formulation
Given a dataset $\mathcal{D}$, the model’s answer for each question $Q_{i}$ in the dataset is answer $A_{i}$ (where $i$ indexes the questions in the dataset). To measure how calibrated the model is, we need three steps. First, we apply an evaluator to get the ground-truth correctness distribution $P_{G_{i}}$ where $G_{i}$ is the random variable that denotes the correctness score in answer $A_{i}$. Formally,
$P_{G_{i}}(x)=\text{Pr}(A_{i}\text{ is }x\%\text{ correct})$ (1)
for $x\in[0,1]$. Second, we leverage a confidence elicitation method to extract confidence distribution $P_{C_{i}}$ from an LLM $\mathcal{M}$, where $C_{i}$ is the random variable that denotes the confidence score of the model $\mathcal{M}$ in answer $A_{i}$. We further assume that the confidence scores are normalized across $x$ such that they form a valid distribution, with the same domain as the correctness distribution. Formally, for $x\in[0,1]$
$P_{C_{i}}(x)$ $=\text{Pr}(\mathcal{M}\text{ thinks }A_{i}\text{ is }x\%\text{ correct})$ (2)
Finally, we use $\mathbb{P_{G}}(\mathcal{D})=\{P_{G_{1}},P_{G_{2}},...,P_{G_{n}}\}$ and $\mathbb{P_{C}}(\mathcal{D})=\{P_{C_{1}},P_{C_{2}},...,P_{C_{n}}\}$ to denote the set of correctness and confidence distributions across the whole dataset, where $n$ is the number of data points. In our framework, a model is calibrated if its confidence distribution for its generated response closely matches the ground-truth correctness distribution, as measured by a metric $M(\mathbb{P_{G}}(\mathcal{D}),\mathbb{P_{C}}(\mathcal{D}))$.
3.2 Groundtruth Correctness
A key challenge in the above formulation lies in obtaining a reliable measure of ground-truth correctness $G(A_{i})$. For short-form QA, correctness can be estimated via simple string comparison metrics Rajpurkar et al. (2016), but for long-form answers, the most reliable evaluation is usually via human annotation. For automatic evaluation, traditional metrics like BLEU and ROUGE are known to have limitations for evaluating long-form QA Liu et al. (2023a), and in our preliminary explorations, we found task-specific metrics like the disambiguation-F1 metric for ASQA Stelmakh et al. (2022) to correlate poorly with the quality of responses generated by closed-book LLMs prompted without reference texts.
Hence, in this paper, we adopt the GPT-4 model to evaluate answers, which has been shown to correlate with human judgments in long-form generation (Liu et al., 2023a). We ask GPT-4 to range an answer from 0 (worst) to 5 (best). The evaluation template includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer. See Appendix E.1 for details. To address potential scoring variability caused by the criteria ambiguity and the inherent uncertainty of LLM, we ask the evaluator to repeatedly evaluate each answer, creating a distribution of scores that more accurately represent the answer’s correctness. Specifically, given an LLM evaluator $\mathcal{E}$, $N$ evaluations from it, and a score $x\in[0,1]$:
$P_{G_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{E}(A_{i})_{j}=x)$ (3)
When N set to 1, the correctness distribution becomes a point-mass distribution, typically observed in tasks with objective ground-truth correctness.
4 Confidence Elicitation
We develop two methods to access the model’s confidence distribution about its own answer. The first is to explicitly ask the model to self-evaluate its own answer. The second is to implicitly assess the model’s confidence in the answer by sampling multiple responses from it and estimating the consistency between them. These methods circumvent the need for logit-based approaches (Jiang et al., 2020), which often have limitations. Specifically, the normalization of logits-derived probabilities often leads to a skewed distribution, as most small values cluster at the lower end, reducing the model’s discriminative ability and complicating interpretation. Additionally, logit-based methods generally imply an overestimation of confidence (Mielke et al., 2022). Finally, the inaccessibility of model logits in many commercial, closed-source platforms restricts the practicality of using logit-based methods in these contexts.
4.1 Self Evaluation
@@
We prompt the model to perform self-evaluation in order to obtain its estimated confidence distribution. The self-evaluation template closely mirrors the template for correctness distribution but omits the reference answer (see Appendix E.3 for details). To get the confidence distribution, we ask the model to repeatedly evaluate the same answer with sampling decoding, with the subsequent scores interpreted as the confidence distribution. Formally, given an LLM $\mathcal{M}$, $N$ self-evaluations from it, and a score $x\in[0,1]$:
$P_{C_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{M}(A_{i})_{j}=x)$ (4)
@@
Such a sampling method provides a more authentic reflection of the model’s internal distribution than logits. By asking the model to assess an answer multiple times, we capture a range of scores that more accurately represent the model’s confidence, which enhances the reliability of the confidence estimation.
4.2 Self Consistency
@@
Another key indicator of model confidence is the consistency among multiple responses a model provides for a given question. Formally, given a primary answer $A_{i}$, other $N$ answers $A_{i}^{1}...A_{i}^{N}$ sampled from LLM $\mathcal{M}$, a metric for measuring the similarity between two answers $Sim(\cdot,\cdot)$, and a score $x\in[0,1]$:
$P_{C_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(Sim(A_{i},A_{i}^{j})=x)$ (5)
Prior research (Kadavath et al., 2022) indicates that such confidence scores correlate to accuracy when the language model is sufficiently large. While it’s straightforward to determine if two short answers are consistent, evaluating them in long-form generation is more complex. To adapt self-consistency to long-form generation, we introduce four different variations to measure answer similarity in long-form content. 1. Naive: An additional LLM assesses the similarity between two responses. 2. Split: The first answer is divided into individual sentences, and another LLM evaluates the presence of similar statements in the second answer. The similarity score is the average of these sentence comparisons. 3. Claim: Building on the Split method, this variation uses a factual detector to identify factual claims in the first answer and checks for their presence in the second response. 4. Named Entity Recoginition (NER): This method further refines the analysis by comparing the overlap of named entities between two responses. For a detailed description of these methodologies and the models used, refer to the Appendix B.
Each of these methods caters to different scenarios and levels of analysis depth, from the broad-strokes assessment of the naive approach to the detail-oriented NER method. The choice of method depends on the specific requirements of the task and the desired balance between comprehensiveness and specificity in the consistency analysis.
5 Calibration Metric
We propose three distinct calibration metrics to capture how well-calibrated the model is from different perspectives. The first metric Correlation examines the relationship between the expected confidence and correctness scores across the entire dataset, serving as an indicator of the model’s proficiency in ranking answers. The second, Wasserstein Similarity, measures the congruence between the confidence and correctness distributions for each instance in the dataset, offering a nuanced view of the model’s ability to mirror the shape and spread of the actual correctness in its confidence distribution. The third metric Selective F1 assesses the model’s ability to judiciously choose when to respond, reflecting the utility of the confidence scores in picking out the best responses across the dataset. These metrics, detailed in the following sections, provide diverse insights into calibration, enriching our understanding of LLM reliability from multiple angles.
5.1 Correlation

### v2

@@ Abstract
To enhance Large Language Models’ (LLMs) reliability, calibration is essential—the model’s confidence scores should align with the likelihood of its responses being correct. However, traditional calibration methods typically rely on a binary true/false assessment of response correctness, unsuitable for long-form generations where an answer can be partially correct. Addressing this gap, we introduce a unified calibration framework, in which both the correctness of the LLMs’ responses and their associated confidence levels are treated as distributions across a range of scores. We develop three metrics for assessing LLM calibration and propose confidence elicitation methods based on self-consistency and self-evaluation. Our experiments demonstrate that larger models don’t necessarily guarantee better calibration, that various calibration metrics complement each other, and that self-consistency methods excel in factoid datasets. We also find that calibration can be enhanced through techniques such as fine-tuning, scaling the temperature. Finally, we illustrate one application of long-form calibration through selective answering in long-form responses, optimizing correctness within a constrained API budget.
@@
Figure 2: Overview of our calibration framework. We prompt an LLM to produce an answer to a specific question, assess the answer’s correctness distribution using an evaluator (task-specific metric/GPT-4 metric/human metric), and determine the model’s confidence distribution through self-evaluation or self-consistency approaches. Finally, we calculate the calibration score by comparing the correctness and confidence distributions against our predefined metrics.
@@ 1 Introduction
Confidence calibration in large language models (LLMs) aims to align the model’s internal confidence with a probabilistic perspective of its answers’ correctness (i.e. quality), enhancing reliability and interpretability for aiding human decision-making Kadavath et al. (2022). People intuitively understand and utilize probabilities Cosmides and Tooby (1996), making this approach crucial for practical applications. Conventional calibration Guo et al. (2017) treats answer correctness as binary (true or false) and seeks to align the model’s confidence with the likelihood of model’s answer being correct, typically stated as: “I am xconfident that this answer is completely correct.” However, the correctness of long-form generation is not always either true or false but can be partially correct (Figure 1).
Therefore, a single confidence score for long-form outputs is ambiguous: it can either imply “I am xconfident that the answer is 100% correct” or “I am 100% confident that the answer is xcorrect.” The former fails to capture the graded notion of long-form answer correctness, while the latter focuses on self-evaluation of correctness, rather than calibration as it overlooks the confidence at specific correctness levels.
Addressing this challenge, we propose to conceptualize the model’s confidence as distribution across scores between $[0,1]$ to capture the nuanced understanding of the model of each correctness level of the long-form answer, corresponding to the statement “I am $x$confident that this answer is $y$correct”. Moreover, we also view the correctness of an LLM’s response as a distribution across scores between $[0,1]$ to capture graded and subjective assessments of long-form generations quality. This subjectivity arises from the multifaceted nature of evaluating long-form outputs, where factors like factuality, coherence, clarity, and comprehensiveness each play a role, potentially introducing variability in judgment Bakker et al. (2022). We can then measure both the classical notion of calibration error, averaged across different correctness levels, as well as new notions of alignment between the correctness and confidence distributions and their utility in selective prediction (§ 3.4). Figure 2 shows an overview of our framework, which consists of three modular components: estimating the target correctness distributions, eliciting confidence distributions from LLMs, and measuring calibration between these distributions.
Our unified framework offers three key advantages. 1. Generalizability: Our framework applies to both long-form and short-form generation tasks by representing correctness and confidence as distributions, regardless of whether the correctness of task is binary, continuous, subjective, or objective. 2. Flexibility: The framework is evaluation-metric agnostic, allowing the integration of any metric or confidence elicitation method, and can adapt as evaluation methods evolve. 3. Interpretability: It provides a nuanced view of uncertainty, enabling decision-makers to assess confidence across multiple correctness levels, fostering greater transparency and trust in the model’s outputs.
@@
We leverage our framework to measure calibration for several LLMs on multiple datasets across three long-form QA—ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022)—and one summarization task, CNNDM (Nallapati et al., 2016). Our results show that our methods excel over baselines by leveraging the model’s nuanced confidence distribution, stronger LLMs like GPT-3.5 don’t necessarily guarantee better calibration, that various calibration metrics complement each other, and LLMs exhibit better calibration on factoid datasets than more open-ended datasets. Furthermore, our analysis highlights that fine-tuning and temperature scaling could enhance calibration. Finally, we illustrate a practical application of long-form calibration: employing a cascading strategy (Chen et al., 2023a) for selective answering to optimize the cost-effectiveness of long-form text generation. In this approach, an open-source model initially handles queries and, based on its confidence levels—assessed using our system—a more advanced API model is engaged as needed. This method ensures cost efficiency while maintaining high-performance levels.
In summary, our contributions are:
@@
  • A universal calibration framework for text generation tasks, enhancing LLM evaluation for critical applications.
@@
  • Evidence that calibration can be improved by model fine-tuning and temperature scaling.
  • A cost-effective model usage strategy, illustrating the practicality of long-form calibration in optimizing LLM deployment.
@@ 2 Related Work
Measuring Calibration Calibration (Guo et al., 2017, Minderer et al., 2021)has been widely studied in language models, whose probabilities derived from logits are generally found to not be calibrated (Jiang et al., 2020, Kadavath et al., 2022, Chen et al., 2023d). Standard metrics to measure the calibration include Expected Calibration Error (ECE) for confidence-accuracy disparity (Naeini et al., 2015), Brier Score for mean squared prediction-outcome differences, and AUROC for assessing confidence-based correct answer identification (Boyd et al., 2013, Kuhn et al., 2023). Selective Accuracy@Coverage measures accuracy within the model’s most confident predictions (Liang et al., 2023, Cole et al., 2023). However, these metrics, rooted in a binary notion of correctness, fall short for long-form tasks where correctness spans a range, suggesting a distribution-based approach is more apt.
Improving Calibration Traditional calibration methods focus on post-processing logits (Guo et al., 2017), but with LLMs generating unbounded text, logits could fall short. Thus, extracting better confidence scores (i.e., confidence elicitation) has become crucial for improving calibration. These include: verbalization, which directly asks the model to output its confidence (Lin et al., 2022), consistency, which uses the uniformity of multiple responses to gauge confidence (Kadavath et al., 2022, Kuhn et al., 2023, Cole et al., 2023, Chen et al., 2023c, Tian et al., 2023a, Lin et al., 2023), and the hybrid of both (Xiong et al., 2023, Tian et al., 2023b, Chen and Mueller, 2023). However, these methods often presume binary answer correctness, offering a singular confidence score that fails to capture the nuanced correctness required for long-form tasks. Recent work (Zhang et al., 2024) on long-form generation addresses continuous correctness scores but focuses on aligning an uncertainty score with correctness, rather than on improving calibration.
3 Long-form Generations Calibration
This section formalizes the long-form generation calibration problem (§ 3.1), and introduces three core components in our calibration framework (Figure 2): the correctness distributions of the answers (§ 3.2), the confidence distributions of the LLM on its answers (§ 3.3), and the calibration metrics to measure how well these two align (§ 3.4).
@@ 3.1 Formulation
Given a dataset $\mathcal{D}$, the model’s answer for each question $Q_{i}$ in the dataset is answer $A_{i}$ (where $i$ indexes the questions in the dataset). To measure how calibrated the model is, we need three steps. First, we apply an evaluator to get the target correctness distribution $P_{T_{i}}$ where $T_{i}$ is the random variable that denotes the correctness score in answer $A_{i}$.
$P_{T_{i}}(x)=\text{Pr}(A_{i}\text{ is }s\text{ correct})$ (1)
for $s\in\mathcal{S}$, where $\mathcal{S}$ is the space of correctness levels (e.g., normalized from ordinal scores ranging from 0 to 5). It should be noted that $s$ could theoretically be a continuous value in the range $[0\%,100\%]$. However, since humans tend to make more accurate judgments using discrete ratings due to “rounding bias” (Honda et al., 2022), we approximate continuous correctness with ordinal scores in practice. Second, We use a confidence elicitation method to derive the confidence distribution $P_{C_{i}}$ from LLM $\mathcal{M}$, where $C_{i}$ represents the model’s confidence in its answer $A_{i}$. We ensure these confidence scores are normalized to form a valid distribution, matching the domain of the correctness distribution. For $s\in\mathcal{S}$,
$P_{C_{i}}(x)=\text{$\mathcal{M}$'s confidence that {A} is {s} correct}$ (2)
Finally, we design metrics to measure the alignment between $P_{C_{i}}$ and $P_{T_{i}}$ across the dataset.
3.2 Correctness Distribution Estimation
To establish correctness distributions as alignment targets, we need to adopt long-form evaluation metrics that integrate aspects such as relevance, coherence, factuality, and helpfulness. Traditional metrics like BLEU and ROUGE fail to capture semantic meaning Liu et al. (2023a), while factuality-based metrics like FactScore Min et al. (2023) may neglect question relevance. GPT-4 metrics have gained popularity Li et al. (2024) due to their adaptability and comparative accuracy (Jain et al., 2023, Liu et al., 2023a). These metrics allow for the integration of various user-prioritized aspects by adjusting evaluation rubrics, providing a balanced approach to both referenced and divergent answers. However, they also have limitations like a bias toward longer outputs Zheng et al. (2023).
Our framework is evaluator-agnostic, allowing us to use any correctness distribution metric. To identify a practical and actionable metric that best guides human decision-making, we conduct human evaluations to determine alignment with human judgment. According to the results in Appendix C, we determined that the GPT-4 metric, with a higher correlation to human judgments compared to the task-specific metric, is more effective for three datasets (e.g., 76.2 v.s 47.8 in ASQA), while the task-specific metric is preferred for another dataset. Our framework’s modular design allows replacements of evaluation metrics based on user needs or new developments.
3.3 Confidence Distribution Elicitation
There are two common strategies to develop confidence in model responses: explicitly asking the model to verbalize its confidence or implicitly estimating it through self-consistency. However, the single confidence score provided by prior studies is ambiguous and non-interpretable in long-form calibration. Therefore, we develop two methods tailored for long-form calibration accordingly.
@@ 3.3 Confidence Distribution Elicitation
Continuous Self Evaluation (CSE) We prompt the model to repeatedly perform self-evaluations, where the resulting scores (typically ordinal scores from 0 to 5) are normalized to a [0,1] range and interpreted as a confidence distribution. The self-evaluation template closely mirrors the template for correctness evaluation but omits the reference answer (see Appendix F.3 for details). Formally, given an LLM $\mathcal{M}$, $N$ self-evaluations:
$P_{C_{i}}(s)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{M}(A_{i})_{j}=s)$ (3)
@@
for score $s\in\mathcal{S}$, where $\mathcal{S}$ is the space of correctness levels. Such a sampling method provides a more authentic reflection of the model’s internal distribution than logits (Cole et al., 2023). By asking the model to assess an answer multiple times, we capture a range of scores that better represent the model’s confidence, which enhances the reliability of the confidence estimation.
@@
Pairwise Self Consistency (PSC) Another key indicator of model confidence is the consistency among multiple responses a model provides for a given question. Given a primary answer $A_{i}$, other $N$ answers $A_{i}^{1}...A_{i}^{N}$ sampled from LLM $\mathcal{M}$, a metric for measuring the similarity between two answers $Sim(\cdot,\cdot)$, and a score $s\in\mathcal{S}$:
$P_{C_{i}}(s)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(Sim(A_{i},A_{i}^{j})=s)$ (4)
Assessing similarity in long-form answers is more complex than with short responses (Kadavath et al., 2022). To address this, we propose four methods for evaluating similarity in long-form content: 1. Naive, assessing overall response similarity with an LLM; 2. Split, analyzing sentence-level similarity; 3. Claim, evaluating claim matching; 4. Named Entity Recognition (NER), focusing on named entity overlap. These approaches range from broad to detailed analysis, chosen based on task requirements and the desired analysis depth. See Appendix B for more details.
3.4 Calibration Metrics
We introduce three key metrics to assess model calibration from various angles. Expected correctness error with multi-class (ECE-M) measures the alignment between the model’s stated confidence in reaching a particular level of correctness and the actual likelihood that the model performs at that specified level, across the spectrum from 0 to 1. Correlation evaluates the alignment between expected confidence and correctness across the dataset, indicating the model’s proficiency in ranking answers. Selective F1 measures the utility of confidence scores in identifying the good answers and abstaining from the rest.
ECE-M The classical notion of calibration relies on an (answer-correctness) pair of random variables $(A,Y)$ $\in\mathcal{A}\times\{0,1\}$, where $\mathcal{A}$ is the answer space. An LLM $\mathcal{M}$ with its confidence elicitation method $h_{\mathcal{M}}$: $\mathcal{A}\rightarrow[0,1]$ is said to be well-calibrated if $\text{Pr}(Y=1|h_{\mathcal{M}}(A)=q)=q$ for $q\in[0,1]$. To measure if this holds, traditional ECE(h) Gupta and Ramdas (2021) is: $\mathbb{E}_{A}\left[\left|\text{Pr}(Y=1|h_{\mathcal{M}}(A))-h_{\mathcal{M}}(A)\right|\right]$
In long-form calibration where the answer correctness is a continuum $Y\in[0,1]$, the probabilistic confidence predictor $h_{\mathcal{M}}$ should predict confidence about each level of correctness and therefore denoted as $h_{\mathcal{M}}$: $\mathcal{A}\times[0,1]\rightarrow[0,1]$. The notion of long-form calibration is $\text{Pr}(Y=s|h_{\mathcal{M}}(A,s)=q_{s})=q_{s}$ for every $s\in[0,1]$ and $q_{s}\in[0,1]$ We define ECE-M as the aggregation of ECE scores for all correctness levels. In practice, we use discrete levels $s\in\mathcal{S}$ (e.g., ratings from 0-5) for the correctness scores. Hence, we calculate an ECE$(s,h)$ conditioned on each $s$,
$\mathbb{E}_{X}\left[\left|\text{Pr}(Y=s|h_{\mathcal{M}}(A,s))-h_{\mathcal{M}}(A,s)\right|\right]$ (5)
Then the final ECE-M score is weighted by the frequency of each class:
[…]